OpenStudy (anonymous):

Solve for the variable. y + -4 = -15

OpenStudy (kaiz122):

add 4 both sides y-4+4=-15+4 y=-11

OpenStudy (anonymous):

y = -15 + 4 y = -11 is the answer...